Question 35
ProbabilityThree students are working independently on a Mathematics problem. Their respective probabilities of solving the problem are 0.6, 0.7 and 0.8. What is the probability that at least one of them solves the problem?
- A.0.024
- B.0.336
- C.0.664
- D.0.976
Show the answer and explanation
Answer: D (0.976)
Explanation
P(none solves) = 0.4 × 0.3 × 0.2 = 0.024, so P(at least one) = 1 − 0.024 = 0.976.